### v1

@@ 4 Internal kinematics
The particular spatial distribution of the MUSE pointings and the limited area they cover prevent us from finding . Therefore, we fixed the central coordinates and of the kinematic field to be either that of the stellar component or that of the HI gas (see values in Table 1). The free parameters were then the systemic l.o.s. velocity , the l.o.s. velocity dispersion common to both models, and the linear gradient k. We assumed the following prior ranges: , where is the mean value of the input velocity distribution, , . The analysis also returned the model evidence Z which resulted useful for comparing the statistical significance of one model against another through the use of the Bayes factor . We obtained the kinematic parameters and the membership probabilities using the MultiNest code (Feroz et al. 2009; Buchner et al. 2014), a multi-modal nested sampling algorithm.
@@
Results of the probability-weighted analysis are reported in Table 3. Of the initial 751 targets, 746 have a . We found that rotation is strongly favoured over a dispersion-only model, with a higher evidence for the case that assumes the central coordinates and orientation of the kinematic major axis from the HI. In this case, the systemic l.o.s. heliocentric velocity of the model including a rotation term is km s, the velocity dispersion km s, and the velocity gradient was km s arcmin (see Fig. 7). As these values have the highest evidence, we adopt these as our reference values in the following analysis.
@@
The systemic velocity and velocity dispersion values we found are within 1- from those previously published by Kirby et al. (2014, i.e. km s and km s, respectively). On the other hand, Wheeler et al. (2017), analyzing the data by Kirby et al. (2014), reported a rotation amplitude for IC 1613 which was largely unconstrained and consistent with null. This is not surprising as their data are spatially distributed roughly perpendicularly to the direction of our MUSE data-set, for which we do detect a clear rotation signal. Therefore, this is the first time that rotation of the stellar component is detected with high significance in IC 1613.
@@
We note that the five low probability targets could still belong to IC 1613 based on their spectral type. Only one has , but we checked that the velocity determination in this case was affected by the presence of residual emission lines from the surrounding ionised gas. By avoiding these lines, we recalculated its velocity, which was compatible with , although with too large an error ( km s). The other stars have . Three of these are low S/N young stars with large velocity errors. Interestingly, the remaining star has and km s, which puts it more than 3 away from the expectation of the best fitting rotation model. However, it is found in F1 and has photometric properties and metallicity (see Sect. 5) in common with the RGB stars of the galaxy. Given the low probability of being a MW contaminant, this star could be a possible binary, but we cannot exclude that it could belong in projection to an extended hot stellar halo around IC 1613. Indeed, Pucha et al. (2019) have shown that its evolved stellar population extends to , while there is growing evidence that MW satellites have extended stellar halos (e.g. Chiti et al. 2021; Longeard et al. 2023; Sestito et al. 2023a, b; Jensen et al. 2024).
@@
4.1 Kinematic properties along each pointing
| Sample | N | |||
| (km s) | (km s) | |||
@@ 4.1 Kinematic properties along each pointing
Here we determine the kinematic properties of IC 1613’s stellar component as a function of radius by analysing the l.o.s. velocities of probable members within each MUSE pointing independently using a dispersion-only model. Results are reported in Table 4 and shown in Fig. 8, where we also compare them to the kinematic properties of the HI component, derived in Read et al. (2016) from the Little-THINGS survey data (Hunter et al. 2012; Oh et al. 2015). To make a direct comparison between gas and stars, we use the same central coordinates, in this case those of the HI, as there is a slight offset with the optical values (see Table 1)151515We note that throughout the text, velocities with the sub-script rot are the observed rotational velocities, neither corrected for the inclination nor for the asymmetric drift; those with the sub-script circ are circular velocities, thereby corrected for both effects.. We verified that the bias introduced by the rotation in the recovery of the velocity dispersion values in each field is negligible in our case (see also Leaman et al. 2012).
@@
It is evident that the typical velocities of stars in each field change with respect to the systemic velocity when moving outward from the galaxy’s centre (see the left panels of Fig. 8, where we define the rotational velocity as as in order to show a positive trend). F1 and F2 contain almost 90% of the inspected stars and thus dominate the linear rotation signal recovered in the analysis of the entire sample (shown as a black solid line in figure). Indeed, the velocity difference between them is km s which, considering that the average distance between F1 and F2 is arcmin, is in agreement with the observed velocity gradient of km s arcmin.
@@
From Fig. 8, we see that F2 and F3 show approximately the same rotational velocity (since their km s). Due to the limited spatial sampling, it is difficult to determine whether the velocity signal has reached a constant value or whether it will continue to vary radially. We recall that the only available spectroscopic dataset that we could use to improve the spatial sampling is that published by Kirby et al. (2014) which, unluckily, is mainly distributed along the optical minor axis of the galaxy providing, as already shown by Wheeler et al. (2017), a poor constraint to the rotation signal. A similar behaviour at these radii is also seen in the HI component, and it appears to be only a local flattening of the rotation curve, as it becomes clear when looking at the HI kinematics over a larger radial range (see also Oh et al. 2015; Read et al. 2016).
@@
Inspection of the velocity dispersion of the stellar component reveals a radial decrease, with the of the central pointing being close to the value obtained from the full sample, while for F2 and F3 we find a significant decrease to a roughly similar lower value. The velocity dispersion profile of the HI starts at km s near the centre of the galaxy, then rises up to km s at (i.e., around the position of F2), to finally decrease and stabilise around km s for . The velocity dispersion of the HI is systematically lower than that obtained for F2 and F3 at the same radii, while this difference increases significantly when compared to the F1 value. It is probable that the lower central HI density due to the presence of a shell-like gas structure could have had an impact on the gas kinematics, and hence on the velocity dispersion profile, inside (see Moiseev & Lozinskaya 2012; Stilp et al. 2013). The HI velocity dispersion is indeed correlated to the gas surface mass density, which regulates the amount of local turbulence (e.g. Bacchini et al. 2020). We will investigate further in the future whether this may be related to the greater velocity dispersion measured from the stars in F1.
@@
In order to estimate the galaxy’s circular velocity and directly compare the kinematics of stars and gas, it was necessary to remove the random motion component that suppresses the rotation curve. We followed the formalism of Read et al. (2016, using their Eqs. 9 and 12) to apply the so-called asymmetric drift correction. Although their analysis focused on the HI component, for our purposes the same reasoning could be applied to the stellar field. We also assumed the same inclination angle (see Table 1) for both tracers. After applying this correction, the stellar and HI components showed excellent agreement. This indicates that the differences observed in the rotational velocities and velocity dispersion between the HI and stellar components are indeed real.
We now examine the ratio of rotational to dispersion support for both stars and gas in IC 1613. For the stellar sample, we obtain a at the projected half-light radius , depending on whether we adopt as the average of the values between F2 and F3 or whether we extrapolate the values obtained for the linear rotation model reported in Table 3. We can consider the values from these two approaches as compatible with each other, taking into account the errors, as shown in Fig. 8. This is the observed without taking into account the inclination angle of IC 1613.
Using low-resolution HI data, Lake & Skillman (1989) reported a value of , in very good agreement with the best-fitting value from Read et al. (2016, see Table 1), which we have adopted. A similar value was also obtained from Oh et al. (2015), although unconstrained. However, Read et al. (2016) suggested that the true inclination could be as low as , allowing them for a better dynamical modelling. This is in contrast to Lake & Skillman (1989), who argue for an inclination angle of . This conclusion is based on the fact that the external isophotes of the HI agree well in shape and orientation with those of the stellar component, as is also the case between the spatial and kinematic P.A. of the HI. If the galaxy was seen almost face-on, it would be difficult to explain such features.
Depending on the inclination adopted, the stellar is between 0.8 and 1.1 for or between 1.9 and 2.7 for . The HI, for comparison, has an observed at the same radius, which rises to when the inclination angles are taken into account.
We can compare our results with those of two similarly bright LG dwarf irregular galaxies, WLM and NGC 6822. Unlike IC 1613, these are seen almost edge-on. WLM shows a stellar at , with the rotation axis aligned with that of the HI (Leaman et al. 2012). In contrast, the kinematics of NGC 6822 is more complex, as is the spatial distribution of its tracers. In fact, the P. A. of the HI is roughly perpendicular to the optical, and while both the stellar and HI components show signs of rotation (Weldrake et al. 2003; Demers et al. 2006; Swan et al. 2016; Wheeler et al. 2017; Belland et al. 2020), it is not clear whether they share the same kinematic axis. In any case, the stellar rotation signal seems to be between 0.4 and 1 (Swan et al. 2016; Wheeler et al. 2017; Belland et al. 2020). For both systems, the HI rotation resulted in much higher values with within for NGC 6822 (Weldrake et al. 2003), and for WLM (Oh et al. 2015; Iorio et al. 2017), similar to the IC 1613 case (assuming though the lowest inclination angle). Our results seem to be in good agreement with the conclusions drawn by Wheeler et al. (2017) that the gas-rich systems of the LG are not strong rotators in their stellar components (i.e. they have ).
@@ 4.2 Kinematic properties as a function of stellar type
We now divide our main sample of probable member stars into two sub-samples, young MS stars and evolved red giant stars, to examine the kinematic properties as a function of stellar type. Relying on the analysis in Sect. 3.1, we perform a simple selection according to the measured : red giant stars (RGS, i.e., belonging to the RGB and AGB with stellar ages Gyr) as those with K, and MS stars (having Gyr) as those with K. We verified that the limit is in general effective in separating the two samples, comparing with the expected from the colours of our targets using the empirical calibration for red giant stars from Alonso et al. (1999), valid in the colour range of and mostly metallicity independent.
@@
The results are reported in Table 4 and Fig. 8. The results for the RGS are very similar (within 1-) to those of the full sample in terms of rotation velocities and velocity dispersion per field, as well as the global values of the linear velocity gradient. This is expected, as this sub-sample represents a large fraction of the main one. As is evident from Fig. 8, a model allowing for a flat rotation curve, or a more steeply rising rotation curve than the one we adopted, would have been a much better match to the kinematics of the MS stars. At all radii, the MS stars are showing a higher amplitude in their rotation with respect to the RGS (and the full sample), and in F2 and F3 also a significantly lower velocity dispersion. From the same figure, we can see that the MS stars follow the HI kinematic properties very closely (apart from the velocity dispersion in F1), while on the other hand, the kinematics of the older RGS, have decoupled from those of the HI. Therefore, the kinematics of the young MS stars appear to be still coupled to that of the HI component.
We checked for the presence of possible biases that could alter the recovery of the kinematic properties of the MS stars. On one hand, the region approximately separating the MS from the RGS, (see Fig. 4), contains the instability strip of IC 1613, where it is expected to find mainly classical Cepheids with short-term variability (i.e., with log and Gyr; Udalski et al. 2001; Bernard et al. 2010, see e.g.). Since we observe them in random phases, we expect any differences to be averaged out. It is also true that, especially for Cepheids, hydrogen or metallic lines can lead to different velocities (see, e.g., Vinko et al. 1998). Nonetheless, we find no significant differences when calculating the average velocities and velocity dispersion including or excluding stars in that color range. Similar considerations can be made for the Be-star candidates in our sample, which should also be short-period variables (e.g., Porter & Rivinius 2003; Rivinius et al. 2013), although in this case it is mostly the intensity of the emission lines that is affected. We repeated the same test finding values compatible within the errors. As for the RGS sample, this contains a small number of C-star candidates, which are long period variables (e.g., Menzies et al. 2015). However, we have verified that the C-stars have similar properties to the rest of the sample within the errors.
We did not check for the presence of binary stars, which could particularly affect the velocity dispersion values. The low S/N of the individual exposures for each field in our sample and the short period over which they were observed (between a few days and a month), preclude a quantitative analysis. We can, however, make qualitative estimates. For a sample of mostly RGB stars with high measured velocity dispersion ( km s), as it is our case, the impact of unresolved binaries is expected to be minimal (i.e. of inflation; Minor et al. 2010; Minor 2013; Spencer et al. 2017, 2018; Arroyo-Polonio et al. 2023, see). However, the case is different when focusing on the MS stars sample. The multiplicity fraction (i.e. the fraction of stars having at least one companion) is generally expected to be for A-types, which increase to for OB stars (e.g. Moe & Di Stefano 2017; Offner et al. 2023). In particular, these stars are often found in triples or higher order multiples. For Be stars such figures are less constrained, but mass accretion through binary interactions has been proposed as a possible channel for their formation (Pols et al. 1991; de Mink et al. 2013; Shao & Li 2014; Bodensteiner et al. 2020).
Considering that the multiple OBs cover a range of orbital periods, preferably a few days or a few decades (Moe & Di Stefano 2017), the stacking process to generate the spectra should already average the short-period ones. For the long-period ones (i.e. log) we expect velocity variations of km s, which could translate into a velocity dispersion inflation of (following Minor et al. 2010, assuming that they behave as short-period low-mass binaries and have intrinsic km s). This would especially affect the central pointing, where the fraction of OB stars is high (). It could thus partly explain why in this field the value of the velocity dispersion of the MS stars is comparable to that of the RGS, while in the other fields it tends to be closer to that of the HI. Another possible source of disturbance for the MS stars in F1 is that they may still be coupled to the motion of the ionised bubble, which is probably expanding (with a typical velocity of km s, Pokhrel et al. 2020). We also do not exclude the presence of possible runaway OB stars that could further impact the kinematics of the sample (e.g. Stoop et al. 2024).
In summary, the different examined tracers point to a similar picture: both IC 1613’s gaseous and stellar components are rotating, with the older stars decoupled from the HI motion mainly by the effect of asymmetric drift, while on timescales of Gyr, the stars have maintained a similar kinematics as the neutral gas. The velocity dispersion profile decreases with radius for all tracers, although in F1 the value of the MS stars may be affected by the presence of binaries or by the expanding motion of the central hole, which may have also affected the central HI velocity dispersion. Nevertheless, there seems to be a correlation between the velocity dispersion of the different tracers and their age, with values increasing as we move towards the older stars. This is in qualitative agreement with the results of Leaman et al. (2017), suggesting that stars are born from gas that becomes less turbulent over time due to decreasing gas fractions.
4.3 Dynamical mass estimation
We can use the results from the previous sections to estimate IC 1613’s dynamical mass, based on the stellar rotational velocities, corrected for the galaxy’s inclination, and the asymmetric drift; therefore turned into circular velocities. We perform this estimate within the 3D de-projected half-light radius to make a direct comparison with values from mass estimators for dispersion supported systems.
We took the average of the observed values of and between F2 and F3, assuming that both the rotation and dispersion profiles are rather flat up to the half-light radius. Assuming both values quoted in Read et al. (2016) for the the galaxy’s inclination i, we would have km s and km s, respectively. The dynamical mass is then between and , respectively. The corresponding mass-to-light ratio within the half-light radius would be between and .
For comparison, the Wolf et al. (2010) mass-estimator for dispersion-supported systems, that is , where G is the gravitational constant, yields within the same radius. This is within the previous results accounting for inclination, but also with the value reported by Kirby et al. (2014). We note that this mass estimator is valid as long as the system is spherically symmetric, without rotation and with a flat l.o.s. velocity dispersion profile. To some degree, all these conditions are violated for IC 1613.
It is important to note that the estimates of the dynamical mass, and dynamical mass-to-light ratio, based on rotational velocities given in this section should be considered as lower limits. The HI circular velocity curve of IC 1613 is known to rise very slowly, even with respect to the expectations of a cored dark matter density profile and assuming the more favourable inclination of . This appears to be the consequence of depressions, particularly evident around 0.5 and 1.5 kpc, due to the presence of large HI bubbles, as discussed in Read et al. (2016). If the stellar velocities reflect a similar behaviour, this would imply that also the stellar rotation curve suffers from these dips, thereby leading to an underestimate of the circular velocity and of the dynamical mass.

### v2

@@ 4.1 Kinematic properties of the main sample: detection of a rotation signal
For the kinematic analysis of IC 1613, the free parameters to be fitted were the systemic l.o.s. velocity , the l.o.s. velocity dispersion , and the linear gradient k for the rotation model. In the latter case, the particular spatial distribution of the MUSE pointings and the limited area they cover prevent us from finding . Therefore, we fixed the central coordinates and of the kinematic field to be either that of the stellar component or that of the HI gas (see values in Table 1).
We assumed the following prior ranges: , where is the mean value of the input velocity distribution, , . The analysis also returned the model evidence Z which resulted useful for comparing the statistical significance of one model against another through the use of the Bayes factor . We obtained the kinematic parameters and the membership probabilities using the MultiNest code (Feroz et al. 2009; Buchner et al. 2014), a multi-modal nested sampling algorithm.
We used the Besançon model of Galactic foreground stars (Robin et al. 2003) as a prior over the expected l.o.s. velocity distribution of the contaminants. We generated a catalogue along IC 1613’s direction, over an area up to its half-light radius, covering the range of colours and magnitudes of our targets. This resulted in a velocity distribution that was well sampled and representative of the spatial area around the MUSE pointings. We approximated it as the sum of two Gaussian profiles (with means km s-1, km s-1, and standard deviations km s-1, km s-1, with an amplitude ratio of ). According to this model, considering the small area covered by the MUSE pointings and the spanned range of magnitudes and colours, we expect to find, in total, around five foreground contaminants in the velocity range between km s-1 and 200 km s-1. If we restrict to the velocity range covered by the input targets (i.e. between km s-1 and km s-1, see Fig. 6), we would expect only one contaminant. We recall that already during the spectral classification step (see Sect. 3.1.1) we found and removed six likely contaminants, all being cold MS stars with velocities around .
@@
Results of the probability-weighted analysis are reported in Table 3. Of the initial 751 targets, 746 have a . We found that rotation is strongly favoured over a dispersion-only model, with a higher evidence for the case that assumes the central coordinates and orientation of the kinematic major axis from the HI. In this case, the l.o.s. systemic velocity of the model including a rotation term is km s-1, the velocity dispersion km s-1, and the velocity gradient was km s-1 arcmin-1 (see Fig. 7). As these values have the highest evidence, we adopt these as our reference values in the following analysis.
@@
The systemic velocity and velocity dispersion values we found are within 1- from those previously published by Kirby et al. (2014, i.e. km s-1 and km s-1, respectively). In contrast to this previous work, we detect a clear rotation signal. In particular, Wheeler et al. (2017), analysing data from Kirby et al. (2014), reported a rotation amplitude for IC 1613 that was largely unconstrained and consistent with a null value. This is not surprising as these data are spatially distributed along the kinematic minor axis (i.e. roughly perpendicular to the direction of our MUSE dataset). Therefore, this is the first time that stellar rotation is detected with high significance in IC 1613.
@@
We note that the five low probability targets could still belong to IC 1613 based on their spectral type. Only one has , but we checked that the velocity determination in this case was affected by the presence of residual emission lines from the surrounding ionised gas. By avoiding these lines, we recalculated its velocity, which was compatible with , although with too large an error ( km s-1). The other stars have . Three of these are low S/N young stars with large velocity errors. Interestingly, the remaining star has and km s-1, which puts it more than 3 away from the expectation of the best fitting rotation model. However, it is found in F1 and has photometric properties and metallicity (see Sect. 5) in common with the RGB stars of the galaxy. Given the low probability of being a MW contaminant, this star could be a possible binary, but we cannot exclude that it could belong in projection to an extended hot stellar halo around IC 1613. Indeed, there is growing evidence that MW satellites have extended stellar halos (e.g. Chiti et al. 2021; Longeard et al. 2023; Sestito et al. 2023a, b; Jensen et al. 2024). Additionally, Pucha et al. (2019) have shown that the evolved stellar population of IC 1613 extends to .
@@
4.2 Kinematic properties of each pointing
| Sample | NP | |||
| (km s-1) | (km s-1) | |||
@@ 4.2 Kinematic properties of each pointing
Here we determine the kinematic properties of IC 1613’s stellar component as a function of radius by analysing the l.o.s. velocities of probable members within each MUSE pointing independently using a dispersion-only model. Results are reported in Table 4 and shown in Fig. 8, where we also compare them to the kinematic properties of the HI component, derived in Read et al. (2016) from the Little-THINGS survey data (Hunter et al. 2012; Oh et al. 2015). To make a direct comparison between gas and stars, we use the same central coordinates, in this case those of the HI, as there is a slight offset with the optical values (see Table 1)161616We note that throughout the text, velocities with the sub-script rot are the observed rotational velocities, neither corrected for the inclination nor for the asymmetric drift; those with the sub-script circ are circular velocities, thereby corrected for both effects.. We verified that the bias introduced by the rotation in the recovery of the velocity dispersion values in each field is negligible in our case (see also Leaman et al. 2012).
@@
It is evident that the typical velocities of stars in each field change with respect to the systemic velocity when moving outward from the galaxy’s centre (see the left panels of Fig. 8, where we define the rotational velocity as as in order to show a positive trend). F1 and F2 contain almost 90% of the inspected stars and thus dominate the linear rotation signal recovered in the analysis of the entire sample (shown as a black solid line in figure). Indeed, the velocity difference between them is km s-1 which, considering that the average distance between F1 and F2 is arcmin, is in agreement with the observed velocity gradient of km s-1 arcmin-1.
@@
From Fig. 8, we see that F2 and F3 show approximately the same rotational velocity (since their km s-1). Due to the limited spatial sampling, it is difficult to determine whether the stellar velocity signal has reached a constant value or whether it will continue to vary radially. We recall that the only available spectroscopic dataset that we could use to improve the spatial sampling is that published by Kirby et al. (2014) which, unluckily, is mainly distributed along the optical minor axis of the galaxy providing, as already shown by Wheeler et al. (2017), a poor constraint to the rotation signal. A similar behaviour at these radii is also seen in the HI component, where the local depressions of the rotation curve may be due to the presence of large HI holes, particularly evident around 0.5 and 1.5 kpc from the centre, as shown in Fig. 1 and also discussed in Read et al. (2016) and Collins & Read (2022).
@@
Inspection of the velocity dispersion of the stellar component reveals a radial decrease, with the of the central pointing being close to the value obtained from the full sample, while for F2 and F3 we find a significant decrease to a roughly similar lower value. The velocity dispersion profile of the HI starts at km s-1 near the centre of the galaxy, then rises up to km s-1 at (i.e., around the position of F2), to finally decrease and stabilise around km s-1 for . The velocity dispersion of the HI is systematically lower than that obtained for the stars at the same radii, with this difference being much more marked in F1. Again, it is likely that the sudden drop in the central part of the HI velocity dispersion profile (marked with hollow pentagons in Fig. 8) is due to the much lower gas density in the inner region, and thus the presence of a shell-like structure within kpc; Moiseev & Lozinskaya 2012; Stilp et al. 2013, see again Fig. 1, but also). The HI velocity dispersion is indeed correlated to the gas surface mass density, which regulates the amount of local turbulence (e.g. Bacchini et al. 2020).
To corroborate the idea that the sudden drop in velocity dispersion for the HI is related to the HI hole, we also overplotted in Fig. 8 (upper-right panel) an exponential fit to the dispersion profiles of a set of dwarf galaxies with high-resolution HI data having similar mass as IC 1613 (i.e. having km s-1), as modelled in Iorio et al. (2017) and Mancera Piña et al. (2021, 2024). In fact, the dispersion profile is expected to increase exponentially towards the centre, and it seems that this was also the case for IC 1613 before the formation of the HI hole.
@@
In order to estimate the galaxy’s circular velocity and verify that we can directly compare the kinematics of stars and gas, it is necessary to account for the inclination and remove the random motion component that suppresses the rotation curve. We followed the formalism of Read et al. (2016, see further details in Sect. 6.2) to apply the so-called asymmetric drift correction, assuming the same inclination angle (see Table 1) for both tracers. After applying this correction, the stellar and HI components resulted in excellent agreement, as shown in Fig. 8 (upper left panel). We discuss further in Sect. 6.2 the caveats related to the inclination value and their impact on the circular velocity and dynamical mass estimations.
@@ 4.3 Kinematic properties as a function of stellar type
We now divide the main sample of probable member stars into two sub-samples, young MS stars and evolved red giant stars, to examine the kinematic properties as a function of stellar type. Relying on the analysis in Sect. 3.1.1, we perform a simple selection according to the measured : red giant stars (RGS, i.e., belonging to the RGB and AGB) as those with K, and MS stars (including the Be stars) as those with K. We verified that the limit is in general effective in separating the two samples, comparing with the expected from the colours of our targets using the empirical calibration for red giant stars from Alonso et al. (1999), valid in the colour range of and mostly metallicity independent. We note that these two samples are also distinct in age. Based on stellar evolution theory, it is known that the RGS track stars of ages Gyr. While a simple comparison with a set of isochrones (with dex, Girardi et al. 2000; Bressan et al. 2012) tells us that the stars we classify as MS are between .
@@
Results of the kinematic analysis are reported in Table 4 and Fig. 8. The outcomes for the RGS are very similar (within 1-) to those of the full sample in terms of rotation velocities and velocity dispersion per field, as well as the global values of the linear velocity gradient. This is expected, as this sub-sample represents a large fraction () of the main one. As is evident from Fig. 8, a model allowing for a flat rotation curve, or a more steeply rising rotation curve than the one we adopted, would have been a much better match to the kinematics of the MS stars. At all radii, the MS stars are showing a higher amplitude in their rotation with respect to the RGS (and the full sample), and in F2 and F3 also a significantly lower velocity dispersion. From the same figure, we can see that the MS stars follow the HI kinematic properties very closely (apart from the velocity dispersion in F1), while on the other hand, the kinematics of the older RGS, have decoupled from those of the HI. Therefore, the kinematics of the young MS stars appear to be still coupled to that of the HI component, although the large errors in the external pointings weaken this conclusion.
We have verified that the reported kinematic values are generally robust against the presence of possible biases, such as the presence of variable or binary stars, as reported in Appendix B.3. Only for the MS stars in F1, we found that the high fraction of OB stars present could inflate the value of the velocity dispersion towards a higher value. This would partly explain why this value is so close, in absolute terms, to that of the RGS. Nevertheless, it is insufficient to account for the large discrepancy with the HI central velocity dispersion. The simplest explanation for this discrepancy is again related to the presence of the central HI hole, where F1 is located. Extrapolating the HI velocity dispersion from the outer parts towards the centre on the basis of the trend observed in the literature for galaxies of similar mass as IC 1613 (see Fig. 8), we would expect the HI velocity dispersion at the location of the central pointing to be very similar to that measured for the RGS and MS stars. Therefore it is possible that the dispersion and spatial distribution of the MS stars in the central pointing reflects that of the HI gas from which they were born. Such a HI hole could have been caused through stellar feedback mechanisms like a supernova (SN) explosion. The shape and nature of the central ionised shells shown in Fig. 2 are typical of a SN remnant, which seem to support this explanation. In addition, the spatial distribution of the OB stars inside and just outside F1 (Fig. 5) seems to suggest not only a physical association with the ionised shells, but that they may have originated from the compression of cold gas due to SN shock waves (Borissova et al. 2004; Garcia et al. 2009).